OpenStudy (anonymous):

Suppose a runner takes 45 minutes to run a routeat 8mi/h at the beginning of the training season. By the end of the season she can run the same route in 38 min. What is her speed at the end of the season?

OpenStudy (anonymous):

reqd speed =distance / time = (8mi/h *45/60 h) / (38/60 h) = 9.47mi/h (appx.)

OpenStudy (zzr0ck3r):

d = r*t so you have d = 45*8 so you know the distance. no again d = r*t now this time we know distance and time d = 45*8 = r*t = r*38 so we have 45*8 = r*38 so r = (45*8)/38 = 9.47 and this is a rate so 9.47m/hr